A wave is propagated in a waveguide at frequency of 9 GHz and separation is 2 cm between walls find cut off wavelength for dominant mode.


1
Expert's answer
2021-01-10T18:30:20-0500

"f_c=\frac{c}{2a}=\frac{3\cdot 10^8}{2\cdot 2\cdot10^{-2}}=7.5 ~GHz."

"u=\frac{c}{\sqrt{1-(\frac{f_c}{f})^2}}=5.43\cdot 10^8 \frac ms."

"\lambda=\frac uf=\frac{5.43\cdot 10^8}{9\cdot 10^9}=6 ~sm."